A bottle labelled with 12V H2O2 contains 700ml solution.  If a student mixes 300ml water in it what is the gm/ltr strength, normality and volume strength of the final solution?

Grade:Select Grade

1 Answers

Khushhal
one year ago
12 V means 1 L of H2O2
  produces 12 L O2
 . But here the volume of solution is 700 mL.
 
As, Volume strength =5.6× Normality
So, 12V=5.6× normality
 
Normality ( in 1000 mL) = 12/5.6
 =2.14N
 
Now, Normality in 700 mL = 2.14/1000×700=1.5N.
 
Now, as we are adding water in it , so normality will remain same, i.e., 1.5 N.
 
Volume strength =5.6× Normality
Volume strength =5.6×1.5=8.4 V